# Interactive, and Editable Simulation for Autonomous Driving

### v1

@@ Abstract
Simulation frameworks have been key enablers for the development and validation of autonomous driving systems. However, existing methods struggle to comprehensively address the autonomy-oriented requirements of balancing: (i) dynamical fidelity, (ii) photorealistic rendering, (iii) context-relevant scenario orchestration, and (iv) real-time performance. To address these limitations, we present a unified framework for creating and curating high-fidelity digital twins to accelerate advancements in autonomous driving research. Our framework leverages a mix of physics-based and data-driven techniques for developing and simulating digital twins of autonomous vehicles and their operating environments. It is capable of reconstructing real-world scenes and assets (real2sim) with geometric and photorealistic accuracy and infusing them with various physical properties to enable real-time dynamical simulation of the ensuing driving scenarios. Additionally, it also incorporates a large language model (LLM) interface to flexibly edit the driving scenarios online via natural language prompts. We analyze the presented framework in terms of its fidelity, performance, and serviceability. Results indicate that our framework can reconstruct 3D scenes and assets with up to 97% structural similarity, while maintaining frame rates above 60 Hz. We also demonstrate that it can handle natural language prompts to generate diverse driving scenarios with up to 95% repeatability and 85% generalizability.
Index Terms:
Agentic AI, Generative AI, Spatial AI, Scene Reconstruction, Scenario Generation, Autonomous Vehicles, Digital Twins, Real2Sim Transfer
@@ I Introduction
Figure 1: Proposed framework for autonomy-oriented digital twins, which blends photorealism and physical interaction with LLM-driven scenario orchestration for enhanced serviceability. Video: https://youtu.be/ZYlcV2J1MME
@@
The development and deployment of autonomous driving systems requires extensive training/optimization and testing/validation in a wide variety of scenarios to ensure safety, robustness, and scalability. Real-world training/testing is often limited by high costs, large time investments, safety concerns, and the lack of control over creating edge cases or extreme conditions. In this context, digital twins offer a tantalizing alternative by providing controllable, repeatable, and diverse simulations for accelerating synthetic data generation, design optimization as well as systematic verification and validation. However, creating high-fidelity digital twins that simultaneously balance photorealism, physical accuracy, operational flexibility, and real-time performance remains a significant challenge.
@@
To this end, we propose a unified framework (refer Fig. 1) for reconstructing high-fidelity digital twins and reconfiguring them via artificially intelligent (AI) agents. Our framework is designed to meet the specific requirements of autonomy-oriented digital twins: (a) physical accuracy and interactivity to simulate vehicle dynamics, sensor characteristics, and environment physics; (b) geometric and visual fidelity to support feature-rich and reliable perception, (c) contextual understanding for generating diverse driving scenarios, and (d) real-time interfacing with autonomous driving software stacks.
@@
The core novelty of the framework lies in its ability to reconstruct real-world scenes and assets with geometric and photorealistic accuracy and seamlessly blend them with synthetic 3D elements represented by point, surface, or mesh models through a real-time hybrid rendering pipeline. By infusing physical properties into these assets, the framework enables real-time dynamical simulation of driving scenarios, capturing realistic interactions between vehicles, pedestrians, and environmental elements. To further enhance its serviceability, the framework integrates an intuitive human-machine interface (HMI) mediated by large language models (LLMs), which allows users to perform creative and interesting edits to generate a diverse range of driving scenarios in real-time via natural language prompts.
Unlike existing solutions that often focus on isolated aspects like physics (e.g., vehicle dynamics), graphics (e.g., static scene rendering), or context (e.g., predefined scenarios), our framework offers a comprehensive reality-to-simulation (real2sim) workflow with real-time performance. This holistic approach not only enhances simulation fidelity but also enables convenient prototyping of autonomous driving algorithms.
@@
  • High-Fidelity Reconstruction: We provide a method to reconstruct autonomy-oriented scenes and assets from camera and LIDAR data and seamlessly create visually, geometrically, and physically accurate, interactive simulations. This allows data-driven digital twins to co-exist with physics-based digital twins in the same scene.
  • Automated Reconfiguration: We present a hierarchical method comprising a mix of LLM agent(s) as well as rule-based agent(s) to automatically generate driving scenarios based on natural language prompts. The proposed method fosters the creativity of generative AI, while enforcing rule-based bounds to ensure the generation of diverse yet pragmatic scenarios.
@@ II Related Work
Simulation frameworks play a crucial role in the field of autonomous driving, as they alleviate the monetary, safety, spatial, and temporal constraints imposed during physical development and validation. However, state-of-the-art frameworks struggle to deliver a unified solution that combines truly photorealistic rendering with physically accurate, real-time simulations. This includes the system-of-systems level simulation of vehicles (sensors, actuators, vehicle dynamics), environments (structure, appearance, variability), and their complex interactions. Additionally, most of the existing approaches lack efficient user interfaces to orchestrate context-aware scenarios that can capture the diverse nature of real-world driving conditions. Below, we review some of the related works, which can be broadly categorized based on their isolated focus on scene reconstruction, scenario reconfiguration, or autonomy-oriented simulation.
@@ II-A Scene Reconstruction
Figure 2: Proposed approach to (a) photorealistic and geometrically accurate 3D scene reconstruction, (b) hybrid scene composition for physically interactive and graphically realistic simulation, and (c) LLM-guided context-aware scenario reconfiguration.
Scene reconstruction methods vary in complexity, accuracy, and application. Each method has unique strengths, with trade-offs between quality, speed, and computational cost. Manual hand-crafting has been shown to achieve good results in some cases, but it is severely time-consuming, labor-intensive, and requires significant artistic skill. Additionally, manual reconstruction often leaves inaccuracies, which impact the environment representation and sensor simulation. Surface reconstruction methods [2] generate continuous and water-tight 3D surfaces from sparse point clouds, excelling in accuracy, but can struggle with noisy or incomplete data, and generally do not capture the visual appearance. Photogrammetry techniques [3, 4] efficiently synthesize 3D models via images captured from multiple viewpoints by analyzing the overlap and spatial relationships between the images, delivering close geometric and visual resemblance, though they require precise image capture and may struggle with the reconstruction of large-scale assets and scenes. Neural radiance fields (NeRFs) [5, 6] leverage modern deep learning methods to implicitly learn an underlying continuous volumetric scene function using a sparse set of input views to generate 3D photorealistic reconstructions by synthesizing views from different perspectives. This method excels at reconstructing complex geometries and textures but demands high computational resources and is significantly slower in terms of rendering. 3D Gaussian splatting (3DGS) [7, 8] offers a fast and memory-efficient solution for scene reconstruction by employing light-weight 3D Gaussian functions to represent points in space as splats parametrized by their pose, scale, color, and opacity, which collectively contribute to the appearance and geometry of reconstruction. This method often outperforms traditional voxel-based methods in terms of speed and quality, specifically excelling in the reconstruction of large-scale scenes and assets, which are central to autonomous driving simulations. However, there is no “one size fits all” solution to scene/asset reconstruction. Hence, we adopt a mix of some of the above techniques such that they complement each other’s strengths and weaknesses.
@@ II-B Scenario Reconfiguration
Scenario reconfiguration methods offer different approaches to generating or modifying scenarios, each with its strengths and limitations, balancing between customization, automation, flexibility, and computational efficiency. Manual design provides full control over the design process, allowing for highly customized and specific scenarios, but it is time-consuming, labor-intensive, and requires expert skill. Data-driven methods [9] leverage large datasets to automatically generate or modify scenarios, offering scalability and efficiency, but are limited by the quality and diversity of the data used, potentially producing generic or unrealistic results. Adversarial generation techniques [10, 11] use generative models and search algorithms to create or transform scenarios into safety-critical corner or edge cases, but often require significant computational power and are sensitive to sampled data quality. Knowledge-based generation [12] utilizes rules, heuristics, or domain-specific information to generate scenarios, ensuring consistency and relevance, but can be less flexible and may struggle with unexpected or novel scenarios. Overall, most of the aforementioned methods have quite steep learning curves and are very systematic, and therefore limited in terms of creative exploration of novel scenarios. Thus, in our work, we propose a hierarchical method comprising a mix of LLM agent(s) and rule-based agent(s) to generate driving scenarios on demand through user-convenient natural language prompts. The proposed method fosters the creativity of generative AI (LLMs can come up with really interesting suggestions, sometimes even beyond the user’s imagination) while enforcing rule-based bounds to ensure the generation of diverse yet pragmatic scenarios.
@@ II-C Autonomy-Oriented Simulation
Simulation frameworks vary in terms of their focus and fidelity. In the context of autonomous driving, the foci can be broadly categorized into dynamics and graphics. On one end of this spectrum, certain frameworks (e.g., [13, 14]) stress vehicle dynamics or terramechanics simulations from multi-body models to finite element analyses. Such frameworks usually compromise on visual fidelity or real-time performance. On the other end of the spectrum, some frameworks (e.g., [15, 16]) only focus on photorealistic rendering, and claim to achieve high fidelity through real-world data replays or graphical animations within reconstructed scenes. Such methods may not be real-time and are sometimes run offline. Finally, there are some frameworks (e.g., [17, 18]) between these two extremes that try to strike a balance between dynamics and graphics. These methods generally use simplified dynamics representations coupled with physically-based rendering (PBR) to compose driving scenarios with hand-crafted digital assets. Another challenge lies in the fact that such simulated scenarios (static scenes or dynamic assets) are often not representative of any real-world entities, which degrades the simulation fidelity. All in all, none of the existing simulation frameworks supports truly photorealistic rendering coupled with hierarchically consistent dynamics and physical accuracy, while offering extremely intuitive and user-friendly scenario reconfiguration and orchestration. Our research aims to address this issue by proposing an openly accessible, high-fidelity yet real-time framework for democratizing AI-driven digital twins in autonomous driving by fusing literal photorealism and dynamical simulation with LLM-guided scenario editing.
@@ III-A Scene Reconstruction
Figure 3: High-fidelity reconstruction: Digital twin of the CU-ICAR campus showing (a) photorealistic rendering via 3D Gaussian splatting; (b) geometric accuracy w.r.t. 3D LIDAR point cloud data; (c) dynamical realism via 3D surface reconstruction; (d) visualization of reconstructed RGB and depth channels in split halves; (e) co-existence of dynamic 3DGS assets like a passenger car, a traffic cone, and a pedestrian sign with 3DMM assets like a road barrier, cement rubble, and a pedestrian; (f) real-time autonomy-oriented simulation with sensor visualization.
The proposed scene/asset reconstruction pipeline (refer Fig. 2(a-b)) can leverage camera and LIDAR data from hand-held or vehicle-mounted sensors. Here, camera data primarily contributes to photorealism, and LIDAR data ensures geometric precision; both data contribute to hybrid scene composition.
The camera data is first preprocessed to extract frames from continuous video sequences, while ensuring consistent feature matching. We observed that camera resolution, exposure changes, and video frame rate have a significant impact on feature matching, as they directly affect image sharpness, brightness/contrast, and motion blur. The extracted frames are parsed through a structure-from-motion (SfM) pipeline to obtain camera intrinsics, undistorted images, reconstructed poses, and feature points. The feature points are used to initialize 3D Gaussians. The camera intrinsics and reconstructed poses are used to project the 3D Gaussians from those viewpoints and rasterize 2D images. These rasterized images are compared with the corresponding undistorted images to compute the photometric loss. The gradients flow back via a differentiable tile rasterizer to update the 3D Gaussians (pose, scale, color, opacity) by minimizing error between the rasterized output and ground truth images.
Particularly, 3D Gaussian Splatting (3DGS) [7] represents scenes and assets using a finite set of “splats”, denoted as ${\mathcal{G}}=\{\bm{g}\}$, where the intensity distribution of every splat follows a 3-dimensional Gaussian function. Each individual Gaussian, $\bm{g}({o},{\bm{\mu}},{\mathbf{q}},{\bm{s}},{\bm{c}})$ is parametrized by opacity ${o}\in(0,1)$, position ${\bm{\mu}}\in\mathbb{R}^{3}$ represented as Cartesian coordinates, orientation ${\mathbf{q}}\in\mathbb{R}^{4}$ represented as a quaternion, anisotropic scaling factors ${\bm{s}}\in\mathbb{R}^{3}_{+}$, and view-dependent color representation ${\bm{c}}\in\mathbb{R}^{F}$ expressed as coefficients of spherical harmonics. To compute the color $C$ of a rasterized pixel, $\mathcal{N}$ Gaussians that intersect with the pixel are sorted based on their distance to the camera center (sorted by index $i\in\mathcal{N}$) and blended using alpha compositing as:
@@
The LIDAR data is first registered to generate a cohesive point cloud, which is then segmented to filter out any unwanted noise or stray artifacts from the data. Next, the point normals are estimated and aligned using a triangulation surface model. This post-processed (aligned) point cloud data (PCD) is used to reconstruct a 3D mesh model (3DMM) of the scene/asset.
@@
The solution yields an approximation of the indicator function $\widetilde{\chi}$ in a least-squares sense. The reconstructed surface is extracted as an isosurface of this indicator function. PSR ensures smooth, watertight surfaces and is particularly robust to noise and non-uniform sampling, making it well-suited for geometrically reconstructing real-world 3D scenes and assets.
The two reconstructed representations (3DGS and 3DMM) are fused to obtain “literally” photorealistic (refer Fig. 3(a)), geometrically accurate (refer Fig. 3(b) and 3(d)), and physically interactive (refer Fig. 3(c) and 3(e)) 3D scenes/assets. This fusion is enabled by a hybrid rendering pipeline comprising hierarchical passes for 3DGS rendering and physically-based rendering (PBR). Consequently, these reconstructed scenes and assets can co-exist with other (hand-crafted, procedurally generated, etc.) 3DGS or 3DMM assets and scenes (refer Fig. 3(e)), enriching scenario generation capabilities. As an example, Fig 3(f) demonstrates the reconstructed CU-ICAR campus (3DGS) within AutoDRIVE Simulator [19], simulating the autonomy-oriented digital twin of OpenCAV [^2] OpenCAV: https://sites.google.com/view/opencav [20] (3DMM) including vehicle dynamics, sensor physics, and environment interactions.

### v2

@@ Abstract
Simulation frameworks have been key enablers for the development and validation of autonomous driving systems. However, existing methods struggle to comprehensively address the autonomy-oriented requirements of balancing: (i) dynamical fidelity, (ii) photorealistic rendering, (iii) context-relevant scenario orchestration, and (iv) real-time performance. To address these limitations, we present a unified framework for creating and curating high-fidelity digital twins to accelerate advancements in autonomous driving research. Our framework leverages a mix of physics-based and data-driven techniques for developing and simulating digital twins of autonomous vehicles and their operating environments. It is capable of reconstructing real-world scenes and assets with geometric and photorealistic accuracy ($\sim$97% structural similarity) and infusing them with physical properties to enable real-time (>>60 Hz) dynamical simulation of the ensuing driving scenarios. Additionally, it incorporates a large language model (LLM) interface to flexibly edit the driving scenarios online via natural language prompts, with $\sim$85% generalizability and $\sim$95% repeatability. Finally, an optional vision language model (VLM) provides $\sim$80% visual enhancement by blending the hybrid scene composition.
@@ I Introduction
Figure 1: Proposed framework for curating autonomy-oriented digital twins, which blends photorealism and physical interaction with LLM-driven scenario orchestration for enhanced serviceability. Video: https://youtu.be/3ovFiRgbHFc
@@
The development and deployment of autonomous driving systems requires extensive training/optimization and testing/validation in a wide variety of scenarios to ensure safety, robustness, and scalability. Real-world training/testing is often limited by high costs, large time investments, safety concerns, and the lack of control over creating edge cases or extreme conditions. In this context, digital twins offer a tantalizing alternative by providing controllable, repeatable, and diverse simulations for accelerating synthetic data generation, design optimization as well as systematic verification and validation. However, creating high-fidelity digital twins that simultaneously balance photorealism, physical accuracy, operational flexibility, and real-time performance remains a significant challenge – one which has typically been underexplored in the literature.
@@
To this end, we propose a unified framework (refer Fig. 1) for reconstructing high-fidelity digital twins of autonomous vehicles and their operating environments, and reconfiguring them via artificial intelligence (AI) agents. Our framework is designed to meet the specific requirements of autonomy-oriented digital twins: (a) physical accuracy and interactivity to simulate vehicle dynamics, sensor characteristics, and environment physics; (b) geometric and visual fidelity to support feature-rich and reliable exteroceptive perception, (c) contextual understanding for generating diverse and creative driving scenarios, and (d) real-time interfacing with autonomous driving software stacks.
Figure 2: Proposed approach to (a) photorealistic and geometrically accurate 3D scene reconstruction, (b) hybrid scene composition of physics-based and data-driven digital twins for physically interactive and graphically realistic simulation, and (c) LLM-guided context-aware scenario reconfiguration with optional VLM-guided visual enhancement.
@@
The core novelty of our framework lies in its ability to reconstruct real-world scenes and assets with geometric and photorealistic accuracy, and to seamlessly blend them with synthetic 3D elements represented as point, surface, or mesh models, through a real-time hybrid physics and rendering engine. By infusing physical properties into these assets, the framework enables real-time, dynamic simulation of driving scenarios that capture realistic interactions among vehicles, pedestrians, and environmental elements. To further enhance its serviceability, the framework integrates an intuitive human-machine interface (HMI) powered by large language models (LLMs), enabling users to perform creative and interesting edits to generate a diverse range of driving scenarios in real time via natural language prompts. Finally, an optional vision language model (VLM) provides appearance enhancement by blending the hybrid scene composition. Unlike existing solutions that often focus on isolated aspects like physics (e.g., vehicle dynamics), graphics (e.g., static scene rendering), or context (e.g., predefined scenarios), our framework offers a comprehensive reality-to-simulation (real2sim) workflow with real-time performance. This holistic approach not only enhances simulation fidelity but also enables convenient prototyping of autonomous driving systems.
@@
  • High-Fidelity Reconstruction: We propose a multi-modal method to reconstruct real-world scenes/assets from camera and LIDAR data and seamlessly create visually, geometrically, and physically accurate, interactive simulations.
  • Automated Reconfiguration: We present a hierarchical method comprising a mix of LLM-driven and rule-based agent(s) to automatically generate driving scenarios using natural language prompts. The proposed method fosters the creativity of generative AI, while enforcing rule-based bounds to generate diverse yet pragmatic scenarios.
  • Optional Enhancement: We introduce an optional VLM to enhance the raw rendering into a more “realistic” one, while imposing explicit constraints on modifying the pose, scale, or other attributes of the scene/assets.
@@ II Related Work
Simulation frameworks play a crucial role in the field of autonomous driving, as they alleviate the financial, safety, spatial, and temporal constraints imposed during physical development and validation. However, state-of-the-art frameworks struggle to deliver a unified solution that combines truly photorealistic rendering with physically accurate, real-time simulations. This includes the system-of-systems level simulation of vehicles (sensors, actuators, vehicle dynamics), environments (structure, appearance, variability), and their complex interactions. Additionally, most of the existing approaches lack efficient user interfaces to orchestrate context-aware scenarios that capture the diverse nature of real-world driving conditions. Below, we review some of the related works, which can be broadly categorized based on their isolated focus on scene reconstruction, scenario reconfiguration, or autonomy-oriented simulation.
@@ II-A Scene Reconstruction
Figure 3: High-fidelity reconstruction: Digital twin of the CU-ICAR campus showing (a) photorealistic rendering via 3DGS; (b) geometric accuracy w.r.t. 3D LIDAR point cloud data; (c) dynamical interaction via PSR; (d) visualization of reconstructed depth channel; (e) co-existence of 3DGS assets like a passenger car, a traffic cone, and a pedestrian sign with 3DMM assets like a road barrier, cement rubble, and a pedestrian; (f) real-time autonomy-oriented simulation with sensor visualization.
Scene reconstruction methods vary in complexity, accuracy, and application. Each method has unique strengths, with trade-offs between quality, speed, and computational cost. Manual hand-crafting has been shown to yield good results in some cases, but it is severely time-consuming, labor-intensive, and requires significant artistic skill. Additionally, manual reconstruction often introduces inaccuracies that affect the environment representation and sensor simulation. Surface reconstruction methods [2] generate continuous, water-tight 3D surfaces from sparse point clouds with high precision, but can struggle with noisy or incomplete data, and generally do not capture the visual appearance. Photogrammetry techniques [3] efficiently synthesize 3D models via images captured from multiple viewpoints by analyzing the overlap and spatial relationships between the images, delivering close geometric and visual resemblance, though they require precise image capture and may struggle with the reconstruction of large-scale assets and scenes. Neural radiance fields (NeRFs) [4] leverage modern deep learning methods to implicitly learn a continuous volumetric scene function from a sparse set of input views, generating 3D photorealistic reconstructions by synthesizing views from different perspectives. This method excels at reconstructing complex geometries and textures but demands high computational resources and is significantly slower in terms of rendering. 3D Gaussian splatting (3DGS) [5] offers a fast and memory-efficient solution for scene reconstruction by employing lightweight 3D Gaussian functions to represent points in space as splats, parametrized by pose, scale, color, and opacity, which collectively contribute to the appearance and geometry of the reconstruction. This method often outperforms traditional voxel-based methods in terms of speed and quality, particularly excelling at reconstructing large-scale scenes and assets, which are central to autonomous driving simulations. However, there is no “one size fits all” solution to scene/asset reconstruction. Hence, in this work, we propose to couple splat-based representations (for enhanced photorealism) with mesh-based representations (for dynamic interaction) to leverage their strengths and complement their weaknesses.
@@ II-B Scenario Reconfiguration
Scenario reconfiguration methods must strike a balance between customization, automation, flexibility, and computational efficiency. Manual design provides complete control over the design process to create highly customized, specific scenarios, but it is time-consuming, labor-intensive, and requires expert skill. Data-driven methods [6] leverage large datasets to automatically generate or modify scenarios, offering scalability and efficiency, but are limited by the quality and diversity of the data used, potentially resulting in generic or unrealistic outcomes. Adversarial generation techniques [7, 8] use generative models and search algorithms to create or transform scenarios into safety-critical corner or edge cases, but often require significant computational power and are sensitive to the quality of the sampled data. Knowledge-based generation [9] utilizes rules, heuristics, or domain-specific information to generate scenarios, ensuring consistency and relevance, but can be less flexible and may struggle with unexpected scenarios. All in all, most of the aforementioned methods are very systematic and, therefore, limited in terms of creative exploration of novel scenarios. Consequently, recent works [10] have begun exploring LLM-based end-to-end scenario generation, yet many lack real-world grounding to audit realistic bounding and logic. Thus, in our work, we propose a hierarchical agentic AI framework comprising a mix of LLM agent(s) and rule-based agent(s) to generate driving scenarios on demand through user-convenient natural language prompts. The proposed method fosters the creativity of generative AI (LLMs can come up with really interesting suggestions, sometimes even beyond the user’s imagination) while enforcing rule-based checks and bounds to ensure the generation of diverse yet pragmatic scenarios.
TABLE I: Comparison with State-of-the-Art Simulators
Simulator
Photorealistic
Rendering
Accurate
Dynamics
Hybrid Scene
Composition
LLM/VLM
Interface
Real-Time
Performance
Open
Source
CarSim [11]
Chrono [12]
GeoSim [13]
MARS [14]
CARLA [15]
LGSVL [16]
AutoDRIVE (Ours)
indicates complete fulfillment; indicates conditional or partial fulfillment; and indicates non-fulfillment.
@@ II-C Autonomy-Oriented Simulation
Simulation frameworks vary in terms of their focus and fidelity. In the context of autonomous driving, the foci can be broadly categorized into dynamics and graphics. On one end of this spectrum, certain frameworks (e.g., [11, 12]) emphasize vehicle dynamics or terramechanics simulations, ranging from multi-body models to finite- or discrete-element methods. Such frameworks usually compromise on visual fidelity or real-time performance. On the other end of the spectrum, some frameworks (e.g., [13, 14]) only focus on photorealistic rendering, and claim to achieve high fidelity through real-world data replays or graphical animations within reconstructed scenes. Such methods may not be real-time and are sometimes run offline. Finally, there are some frameworks (e.g., [15, 16]) that fall between these two extremes and try to strike a balance between dynamics and graphics. These methods generally use simplified dynamics representations coupled with physically-based rendering (PBR) to compose driving scenarios with hand-crafted digital assets. Another challenge is that such simulated scenarios (static scenes or dynamic assets) are often not representative of real-world entities, which raises questions about simulation fidelity. All in all, none of the existing simulation frameworks support truly photorealistic rendering, high-fidelity dynamical simulation, and real2sim grounding, while offering extremely intuitive and user-friendly scenario reconfiguration and orchestration (refer Table I). Our research aims to address this gap by proposing an openly accessible, high-fidelity yet real-time framework for democratizing AI-driven digital twins in autonomous driving by fusing literal photorealism and dynamical simulation with LLM-guided scenario editing.
@@ III-A Scene Reconstruction
The proposed scene/asset reconstruction pipeline (refer Fig. 2(a-b)) can leverage camera and LIDAR data from hand-held or vehicle-mounted sensors. Here, camera data contributes to photorealism, LIDAR data ensures geometric precision, and their combination enables hybrid scene composition.
The camera data is first preprocessed to extract frames from continuous video sequences, while ensuring consistent feature matching. We observed that camera resolution, exposure changes, and video frame rate significantly impact feature matching by directly affecting image sharpness, brightness/contrast, and motion blur. The extracted frames are parsed through a structure-from-motion (SfM) pipeline to obtain camera intrinsics, undistorted images, reconstructed poses, and feature points. The feature points are used to initialize 3D Gaussians. The camera intrinsics and reconstructed poses are used to project the 3D Gaussians from those viewpoints and rasterize 2D images. These rasterized images are compared with the corresponding undistorted images to compute the photometric loss. The gradients flow back via a differentiable tile rasterizer to update the 3D Gaussians (pose, scale, color, opacity) by minimizing error between the rasterized output and ground truth images.
Particularly, 3D Gaussian Splatting (3DGS) [5] represents scenes and assets using a finite set of “splats”, denoted as ${\mathcal{G}}=\{\bm{g}\}$, where the intensity distribution of every splat follows a 3-dimensional Gaussian function. Each individual Gaussian, $\bm{g}({o},{\bm{\mu}},{\mathbf{q}},{\bm{s}},{\bm{c}})$ is parametrized by opacity ${o}\in[0,1]$, position ${\bm{\mu}}\in\mathbb{R}^{3}$ represented as Cartesian coordinates, orientation ${\mathbf{q}}\in\mathbb{R}^{4}$ represented as a quaternion, anisotropic scaling factors ${\bm{s}}\in\mathbb{R}^{3}_{+}$, and view-dependent color representation ${\bm{c}}\in\mathbb{R}^{F}$ expressed as coefficients of spherical harmonics. To compute the color $C$ of a rasterized pixel, $\mathcal{N}$ Gaussians that intersect with the pixel are sorted based on their distance to the camera center (sorted by index $i\in\mathcal{N}$) and blended using alpha compositing as:
@@
The LIDAR data is first registered to generate a cohesive point cloud, which is then segmented to filter out any unwanted noise and stray artifacts. Next, the point normals are estimated and aligned using a triangulation surface model. This post-processed (aligned) point cloud data (PCD) is used to reconstruct a 3D mesh model (3DMM) of the scene/asset.
@@
The solution yields an approximation of the indicator function $\widetilde{\chi}$ in a least-squares sense. The reconstructed surface is extracted as an isosurface of this indicator function. PSR ensures smooth, watertight surfaces and is particularly robust to noise and non-uniform sampling, making it well-suited for geometric reconstruction of real-world 3D scenes and assets.
Figure 4: Samples of reconstructed scenes (a) CU-ICAR, (b) CGEC, (c) AuE lab, (d) lit footpath, (e) unlit footpath; and assets (f) passenger car, (g) traffic cone, (h) pedestrian sign. Notice the shadows and reflections captured from real-world data.
The hybrid composition of 3DGS and PSR jointly achieves photorealism and physical validity in autonomous driving simulation. 3DGS provides a dense, view-dependent radiance representation with high-fidelity novel view synthesis, but its density-based, non-manifold structure lacks explicit surfaces, leading to unstable depth estimation. PSR, by contrast, reconstructs a globally consistent implicit surface from oriented PCD, yielding watertight meshes with well-defined normals suitable for collision detection and raycasting, but at the cost of limited appearance expressiveness. By decoupling appearance (3DGS) from physics (PSR) with geometrical grounding, the hybrid representation enforces cross-modal consistency across RGB-D channels, resulting in perceptually realistic and physically interactive scenes and assets.
The two reconstructed representations (3DGS and 3DMM) are fused to obtain “literally” photorealistic (refer Fig. 3(a)), geometrically accurate (refer Fig. 3(b) and 3(d)), and physically interactive (refer Fig. 3(c)) 3D scenes/assets. This fusion is enabled by a hybrid rendering pipeline that comprises hierarchical passes for 3DGS rendering and physically based rendering (PBR). Consequently, these reconstructed scenes and assets can co-exist with other 3DGS or 3DMM assets and scenes (refer Fig. 3(e)), enriching scenario generation capabilities. As an example, Fig 3(f) demonstrates the reconstructed CU-ICAR campus (3DGS) within AutoDRIVE Simulator [17], simulating the autonomy-oriented digital twin of OpenCAV [^2] OpenCAV: https://sites.google.com/view/opencav [18] (3DMM) including vehicle dynamics, sensor physics, and environment interactions. Fig. 4 depicts a few examples of reconstructed scenes and assets.